### v1

@@ Abstract
In this paper we present a new “external verifier” for the Lean theorem prover, written in Lean itself. This is the first complete verifier for Lean 4 other than the reference implementation in C++ used by Lean itself, and our new verifier is competitive with the original, running between 20% and 50% slower and usable to verify all of Lean’s mathlib library, forming an additional step in Lean’s aim to self-host the full elaborator and compiler. Moreover, because the verifier is written in a language which admits formal verification, it is possible to state and prove properties about the kernel itself, and we report on some initial steps taken in this direction to formalize the Lean type theory abstractly and show that the kernel correctly implements this theory, to eliminate the possibility of implementation bugs in the kernel and increase the trustworthiness of proofs conducted in it. This work is still ongoing but we plan to use this project to help justify any future changes to the kernel and type theory and ensure unsoundness does not sneak in through either the abstract theory or implementation bugs.
keywords:
@@
Lean, proof assistant, external verifier, implementation, metatheory, type theory, proof theory
category:
1 Introduction
@@
Lean [7] is a theorem prover based on the Calculus of Inductive Constructions (CIC), quite similar to its older brother Coq [3], but as the name suggests, one of the differentiating aspects was the desire to make the kernel “leaner,” relying on simpler primitives while retaining most of the power of the system. This was in particular informed by periodic news of unsoundnesses in Coq [^1] https://github.com/coq/coq/blob/master/dev/doc/critical-bugs.md due to complications of the many interacting features in the kernel. So from the beginning, Lean advertised its simpler foundations and multiple external verifiers as reasons that it would not suffer from the same troubles. These hopes were broadly upheld: for the entire release history of Lean 3, there were no soundness bugs reported against the kernel (written in C++), and our previous work [4] showed the consistency of Lean with respect to ZFC with $n$ inaccessible cardinals for all $n<\omega$, so the soundness story was fairly strong.
@@
But times move on, and Lean 4 now exists as an (almost) ground-up rewrite of Lean with most components now written in Lean itself. The kernel was one of the few components that was not rewritten, but it was extended with various features for performance reasons like bignum arithmetic, nested inductive types, and primitive projections, and unfortunately some soundness bugs crept in during the process and Lean’s record is no longer spotless. Another thing that was lost during the port was the external verifiers: the whole metaprogramming infrastructure was redesigned so old external verifiers are no longer applicable to Lean 4. While most of the new features can be treated as mere abbreviations, nested inductive types and $\eta$ for structures significantly impact the theory, and as a result the soundness proof from [4] is no longer directly applicable. (But see [12], which covers some of these modifications.)
@@
To restore the soundness story of Lean to a satisfactory state, we believe it is necessary to bring back the external verifiers. And what better language to do so than Lean itself? There are reasons not to use Lean for a truly external verifier (and more external verifiers are coming, in Swift [^2] https://github.com/gebner/trepplein and Rust [^3] https://github.com/ammkrn/nanoda_lib), but there are also advantages to having a kernel written in Lean:
@@
In this paper, we will present Lean4Lean, an external verifier for Lean 4, written in Lean. The verifier itself is more than a prototype: it is fully capable of checking the entirety of Lean (the Lean compiler package), Std (the standard library), and Mathlib (the Lean mathematics library). It is not as fast as the original kernel written in C++ because the Lean compiler still has some ways to go to compete with C++ compilers, but the overhead is still reasonable in exchange for the additional guarantees it can provide, and it is suitable as a complementary step in Lean projects that wish to validate correctness in a variety of ways.
The second part of this project, which is much larger and on which we can only report partial progress, is the specification of Lean’s metatheory and verification of the Lean4Lean kernel with respect to that theory. This is most similar to the MetaCoq project [11], and amounts to a formalization of the results of [4]. Our principal contributions in this area is a definition of the typing judgment, along with some regularity theorems we can prove and others we can only state, as well as invariants and proofs of correctness for a few of the components of the kernel. The remainder is left as future work.
@@
2 Base theory
2.1 Expressions
At the heart of the theory is the VExpr [^4] The prefix “V” for “verified” disambiguates it from the Expr type used by the kernel itself. This is the more abstracted version of expressions used for specification purposes. type, which represents expressions of type theory.
@@ 2.1 Expressions
  | bvar (deBruijnIndex : Nat)
  | sort (u : VLevel)
  | const (declName : Name) (us : List VLevel)
  | app (fn arg : VExpr)
  | lam (binderType body : VExpr)
  | forallE (binderType body : VExpr)
@@
  • const \$c\$ \$us\$ represents a reference to global constant $c$ in the environment, instantiated with universes $us$. (Constants in Lean can be universe-polymorphic, and they are instantiated with concrete universes at each use site.)
@@
  • lam \$A\$ \$e\$ is a lambda-expression $\lambda x:A.\,e$. Because we are using de Bruijn variables, $x$ is not represented; instead $e$ is type-checked in an extended context where variable 00 is type $A$ and the other variables are shifted up by 1.
  • forallE \$A\$ \$B\$ [^5] The “E” suffix is used because forall is a keyword; we could use forall anyway but it would require escaping in several places. is a dependent Pi type $\Pi x:A.\,B$, also written as $\forall x:A.\,B$ because for Prop this is the same as the “for all” quantifier. It uses the same binding structure as $\lambda$.
@@
  | zero  : VLevel
  | succ  : VLevel  VLevel
  | max   : VLevel  VLevel  VLevel
  | imax  : VLevel  VLevel  VLevel  imax a b means (if b = 0 then 0 else max a b)
  | param : Nat  VLevel
These come up when type-checking expressions, for example the type of sort u is sort (succ u). The one notable case here is param \$i\$, which represents the $i$th universe variable, where $i<n$ where $n$ is the number of universe parameters to the current declaration.
This definition corresponds to the following BNF style grammar from [4], with the main change being that it is more precise about how variables and binding are represented. (For presentational reasons, we will use named variables in the informal version, but the formalization uses lift and subst functions when moving expressions between contexts.)
$\hskip 110.00017pt\ell$ $::=0\mid S\ell\mid\max(\ell,\ell)\mid\operatorname{imax}(\ell,\ell)\mid u$
@@
2.2 Typing and definitional equality
@@
[…]
Figure 1: The rules for the judgment $\Gamma\vdash e\equiv e^{\prime}:\alpha$, with parameters $E$ (the global environment) and $n$ (the number of universe parameters in context), suppressed in the notation. This is VEnv.IsDefEq in the formalization.
For the typing rules, we use a slight variation of the typing judgments from [4], see Figure 1. Some remarks:
  • In [4], there were two judgments $\Gamma\vdash e:\alpha$ and $\Gamma\vdash e_{1}\equiv e_{2}$ which are mutually inductive (due to t-conv and t-proof-irrel), and $\Gamma\vdash e_{1}\equiv e_{2}:\alpha$ was a defined notion for $\Gamma\vdash e_{1}\equiv e_{2}\ \wedge\ \Gamma\vdash e_{1}:\alpha\ \wedge\ \Gamma\vdash e_{2}:\alpha$. In this version, we have only one all-in-one relation $\Gamma\vdash e_{1}\equiv e_{2}:\alpha$, and we define $\Gamma\vdash e:\alpha$ to mean $\Gamma\vdash e\equiv e:\alpha$ and define $\Gamma\vdash e_{1}\equiv e_{2}$ as $\exists\alpha.\;(\Gamma\vdash e_{1}\equiv e_{2}:\alpha)$.
    We conjecture the two formulations to be equivalent, but this version seems to be easier to prove basic structural properties about (see subsection 2.3). Moreover Lean does not have good support for mutual inductive predicates, so keeping it as a single inductive makes induction proofs easier.
@@
  • The t-extra rule says that we can add arbitrary definitional equalities from the environment. This is how we will add the two supported “extensions” to the theory, for inductive types and quotients, without the details of these extensions complicating reasoning about the core theory. But we can’t actually support arbitrary definitional equalities—instead each theorem about the typing judgment has its own assumptions about what extensions are allowed.
2.3 Properties of the typing judgment
@@
  1. 1.
    (IsDefEq.closedN) If $\Gamma\vdash e\equiv e^{\prime}:\alpha$, then $e$, $e^{\prime}$, and $\alpha$ are well-scoped (all free variables have indices less than $|\Gamma|$).
  2. 2.
@@
  3. 3.
    (IsDefEq.instL) If $\Gamma\vdash_{E,n}e\equiv e^{\prime}:\alpha$ and $\forall i.\;n^{\prime}\vdash\ell_{i}$, then
    $\Gamma[\bar{u}\mapsto\bar{\ell}]\vdash_{E,n^{\prime}}e[\bar{u}\mapsto\bar{\ell}]\equiv e^{\prime}[\bar{u}\mapsto\bar{\ell}]:\alpha[\bar{u}\mapsto\bar{\ell}]$.
  4. 4.
    (IsDefEq.instN) If $\Gamma,x:\beta\vdash e_{1}\equiv e_{2}:\alpha$ and $\Gamma\vdash e_{0}:\beta$, then
    $\Gamma\vdash e_{1}[x\mapsto e_{0}]\equiv e_{2}[x\mapsto e_{0}]:\alpha[x\mapsto e_{0}]$.
Proof 2.2 (Proof sketch).
  1. 1.
    This is a straightforward proof by induction on $\Gamma\vdash e\equiv e^{\prime}:\alpha$, except that one gets stuck at t-const because (assuming the environment is well-typed, see section 4) we know that $\vdash c_{\bar{\ell}}:\alpha$ and need that $\alpha$ is well-scoped. So in fact this is a double induction, first over the environment $E$ and then over $\Gamma\vdash e\equiv e^{\prime}:\alpha$. We also need lemmas about $e[x\mapsto e^{\prime}]$ and $e[\bar{u}\mapsto\bar{\ell}]$ preserving well-scopedness, but these are also direct by induction on $e$.
  2. 2.
    By induction. We use 2.1.1 in the t-const case, to show that the $\alpha$ in $\Gamma,\Gamma^{\prime}\vdash c_{\bar{\ell}}\equiv c_{\bar{\ell}^{\prime}}:\alpha$ can be used in $\Gamma,\Delta,\Gamma^{\prime}\vdash c_{\bar{\ell}}\equiv c_{\bar{\ell}^{\prime}}:\alpha$ without renaming any bound variables because $\alpha$ is closed.
  3. 3.
@@
  4. 4.
    By induction on the first hypothesis. This uses 2.1.2 when lifting $\Gamma\vdash e_{0}:\beta$ under binders.
@@
Proof 2.4.
We have $\Gamma,x:\alpha^{\prime}\vdash x:\alpha$ by t-bvar, t-conv and weakening, and
$\Gamma,\_:\alpha^{\prime},x:\alpha\vdash e_{1}\equiv e_{2}:\beta$ by weakening, so $\Gamma,x:\alpha^{\prime}\vdash e_{1}[x\mapsto x]\equiv e_{2}[x\mapsto x]:\beta[x\mapsto x]$ by 2.1.4.
Lemma 2.5 (IsDefEq.forallE_inv).
If $\Gamma\vdash(\forall x:\alpha.\;\beta):\gamma$ then $\Gamma\vdash\alpha\ \mathsf{type}$ and
$\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$ (and therefore also $\Gamma\vdash(\forall x:\alpha.\;\beta)\ \mathsf{type}$).
Proof 2.6 (Proof sketch).
@@
  • t-all: If $e_{1}$ is $\forall x:\alpha.\;\beta$ we are done; if $e_{2}$ is $\forall x:\alpha.\;\beta$ then we obtain $\Gamma,x:\alpha^{\prime}\vdash\beta\ \mathsf{type}$ and need 2.3 to get $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$.
  • t-beta: For this to apply, it must be that we have $(\lambda y:\delta.\;e)\;e^{\prime}\equiv e[y\mapsto e^{\prime}]$ where $e[y\mapsto e^{\prime}]$ is $\forall x:\alpha.\;\beta$. So either $e=y$ and $e^{\prime}=\forall x:\alpha.\;\beta$, in which case the inductive hypothesis for $e^{\prime}$ applies, or $e=\forall x:\alpha^{\prime}.\;\beta^{\prime}$ with $\alpha^{\prime}[y\mapsto e^{\prime}]=\alpha$ and $\beta^{\prime}[y\mapsto e^{\prime}]=\beta$ in which case $\Gamma,y:\delta\vdash\alpha^{\prime}\ \mathsf{type}$ and $\Gamma,y:\delta,x:\alpha^{\prime}\vdash\beta^{\prime}\ \mathsf{type}$ by the inductive hypothesis for $e$, and 2.1.4 applies.
@@
Lemma 2.7 (IsDefEq.sort_inv).
If $\Gamma\vdash_{E,n}\mathsf{U}_{\ell}:\gamma$ then $n\vdash\ell\ \mathsf{ok}$ (and therefore also $\Gamma\vdash\mathsf{U}_{\ell}:\mathsf{U}_{S\ell}$).
Proof 2.8 (Proof sketch).
Similar to 2.5.
@@
If $\Gamma\vdash e:\alpha$ then $\Gamma\vdash\alpha\ \mathsf{type}$.
Proof 2.10 (Proof sketch).
@@
  • t-app: We have $\Gamma\vdash(\forall x:\alpha.\;\beta)\ \mathsf{type}$ from the IH and $\Gamma\vdash e_{2}:\alpha$ by assumption, and from 2.5 we get $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$, so $\Gamma\vdash\beta[x\mapsto e_{2}]\ \mathsf{type}$ by the substitution lemma.
  • t-all: We have $\Gamma\vdash\mathsf{U}_{\ell_{1}}\ \mathsf{type}$ and $\Gamma,x:\alpha\vdash\mathsf{U}_{\ell_{1}}\ \mathsf{type}$ from the IH, so by 2.7 we have $n\vdash\ell_{1},\ell_{2}\ \mathsf{ok}$, therefore $\Gamma\vdash\mathsf{U}_{\operatorname{imax}(\ell_{1},\ell_{2})}\ \mathsf{type}$.
@@
If $\Gamma,x:\alpha\vdash f\equiv f^{\prime}:\beta$ and $\Gamma,x:\alpha\vdash a\equiv a^{\prime}:\alpha$ then
$\Gamma\vdash f[x\mapsto a]\equiv f^{\prime}[x\mapsto a^{\prime}]:\beta[x\mapsto a]$.
Proof 2.12 (Proof sketch).
@@
$\hskip 100.00015ptf[x\mapsto a]$ $\equiv(\lambda x:\alpha.\;f)\;a$
@@
To finish, we apply the lemma twice, once with $f$ and $\beta$ so that it suffices to show $\beta[x\mapsto a]\equiv\beta[x\mapsto a^{\prime}]:\mathsf{U}_{\ell}$ and then again with $\beta$ in place of $f$ and $\mathsf{U}_{\ell}$ in place of $\beta$ so that it suffices to show $\mathsf{U}_{\ell}[x\mapsto a]\equiv\mathsf{U}_{\ell}[x\mapsto a^{\prime}]:\mathsf{U}_{S\ell}$, which is true by reflexivity.
2.4 Conjectured properties of the typing judgment

### v2

@@ Abstract.
In this paper we present a new “external checker” for the Lean theorem prover, written in Lean itself. This is the first complete typechecker for Lean 4 other than the reference implementation in C++ used by Lean itself, and our new checker is competitive with the original, running between 20% and 50% slower and usable to verify all of Lean’s mathlib library, forming an additional step in Lean’s aim to self-host the full elaborator and compiler. Moreover, because the checker is written in a language which admits formal verification, it is possible to state and prove properties about the kernel itself, and we report on some initial steps taken in this direction to formalize the Lean type theory abstractly and express the relation between the kernel functions and the type theory. We plan to use this project to help justify any future changes to the kernel and type theory and ensure unsoundness does not sneak in through either the abstract theory or implementation bugs.
@@
Lean, proof assistant, external typechecker, implementation, metatheory, type theory, proof theory
ccs: Mathematics of computing Mathematical software performanceccs: Software and its engineering Formal methodsccs: Social and professional topics Systems analysis and design
1. Introduction
@@
Lean (de Moura et al., 2015) is a theorem prover based on the Calculus of Inductive Constructions (CIC) (Paulin-Mohring, 2015), quite similar to its older brother Coq (Team, 2020), but as the name suggests, one of the ways it sought to differentiate itself was the desire to make the kernel “leaner,” relying on simpler primitives while retaining most of the power of the system. This was in particular informed by periodic news of unsoundnesses in Coq [^1] https://github.com/coq/coq/blob/master/dev/doc/critical-bugs.md due to complications of the many interacting features in the kernel. So from the beginning, Lean advertised its simpler foundations and multiple external checkers (Selsam, 2017; Ebner, 2022) as reasons that it would not suffer from the same troubles. These hopes were broadly upheld: for the entire release history of Lean 3, there were no soundness bugs reported against the kernel (written in C++), and Carneiro (Carneiro, 2019) showed the consistency of Lean with respect to ZFC with $n$ inaccessible cardinals for all $n<\omega$, so the soundness story seemed fairly strong.
@@
But times move on, and Lean 4 now exists as an (almost) ground-up rewrite of Lean with most components now written in Lean itself. The kernel was one of the few components that was not rewritten, but it was extended with various features for performance reasons like bignum arithmetic, nested inductive types, and primitive projections, and unfortunately some soundness bugs crept in during the process and Lean’s record is no longer spotless. Another thing that was lost during the port was the external checkers: the whole metaprogramming infrastructure was redesigned so old external checkers are no longer applicable to Lean 4. While most of the new features can be treated as mere abbreviations, nested inductive types and $\eta$ for structures significantly impact the theory, and as a result the soundness proof from (Carneiro, 2019) is no longer directly applicable. (But see (Ullrich, 2023), which covers some of these modifications.)
We interpret this as a cautionary tale: It is not sufficient to see a prior system and hope to do better through sheer force of will. We are all human, and mistakes in both the theory and the implementation happen as a matter of course. The way to do better is not to rewrite, but to set up a process that structurally ensures that mistakes are either prevented or corrected. For a proof assistant, there are three main ways to do this: (1) testing, (2) independent reimplementation, and (3) formal verification. Lean has a fairly robust test suite, but testing helps more in the high-frequency low-impact domain; soundness bugs are rare and generally occur in areas that are insufficiently tested because the developer did not even consider the error condition.
@@
The main contribution of this paper lies in (2): an independent reimplementation. To restore the soundness story of Lean to a satisfactory state, we believe it is necessary to bring back the external checkers. And what better language to do so than Lean itself? There are reasons not to use Lean for a truly independent verifier (and more external checkers are coming, in Swift [^2] https://github.com/gebner/trepplein and Rust [^3] https://github.com/ammkrn/nanoda_lib), but there are also advantages to having a kernel written in Lean:
@@
In this paper, we will present Lean4Lean, an external verifier for Lean 4, written in Lean. The verifier itself is more than a prototype: it is fully capable of checking the entirety of Lean (the Lean compiler package), Batteries (the extended standard library), and Mathlib (the Lean mathematics library). It is not as fast as the original kernel written in C++ because the Lean compiler still has some ways to go to compete with C++ compilers, but the overhead is still reasonable in exchange for the additional guarantees it can provide, and it is suitable as a complementary step in Lean projects that wish to validate correctness in a variety of ways.
The second part of this project, which is much larger and on which we can only report partial progress, is the specification of Lean’s metatheory and verification of the Lean4Lean kernel with respect to that theory, covering mistake prevention measure (3). This is most similar to the MetaCoq project (Sozeau et al., 2020, 2019), and amounts to a formalization of the results of (Carneiro, 2019). Our principal contributions in this area are a definition of the typing judgment, along with some regularity theorems we can prove and others we can only state, as well as some invariants used in the kernel. The remainder is left as future work.
In this paper, we will avoid making value judgments about Lean or Coq and their respective theories. These systems came to be what they are now as a result of complex historical factors, and we seek only to validate the behavior of Lean as it exists, even though it lacks some type theoretic properties that would have made the job easier if true. One possible outcome of this project is a recommendation for how to improve the typing rules, but this comes at a high cost to existing users so we do not take it lightly and we have no such recommendations currently.
@@
2. Base theory
2.1. Expressions
At the heart of the theory is the VExpr [^4] The prefix “V” for “verified” disambiguates it from the Expr type used by the kernel itself. This is the more abstracted version of expressions used for specification purposes. type, which represents expressions of type theory. This is a minimalistic version of the type actually used by the kernel in section 3.1, to ease the burden of proving theoretical properties. [^5] MetaCoq seems not to do this; it uses the same type for proving metatheory and for writing metaprograms (it’s not the same type as the one Coq itself uses because that type is in ML, but it is roughly equivalent).
@@ 2.1. Expressions
| bvar (deBruijnIndex : Nat)
| sort (u : VLevel)
| const (declName : Name) (us : List VLevel)
| app (fn arg : VExpr)
| lam (binderType body : VExpr)
| forallE (binderType body : VExpr)
@@
  • const \$c\$ \$us\$ represents a reference to global constant $c$ in the environment, instantiated with universes $us$. (Constants in Lean can be universe-polymorphic, and they are instantiated with concrete universes at each use site.) Inductive constructors, recursors, and definitions are all represented using this constructor.
@@
  • lam \$A\$ \$e\$ is a lambda-abstraction $\lambda x:A.\,e$. Because we are using de Bruijn variables, $x$ is not represented; instead $e$ is typechecked in an extended context where variable 00 is type $A$ and the other variables are shifted up by 1.
  • forallE \$A\$ \$B\$  [^6] The “E” suffix is used because forall is a keyword; we could use forall anyway but it would require escaping in several places. is a dependent $\Pi$-type $\Pi x:A.\,B$, also written as $\forall x:A.\,B$ because for Prop this is the same as the “for all” quantifier. It uses the same binding structure as $\lambda$.
@@
| zero : VLevel
| succ : VLevel VLevel
| max : VLevel VLevel VLevel
| imax : VLevel VLevel VLevel
imax a b means (if b = 0 then 0 else max a b)
| param : Nat VLevel
These come up when typechecking expressions, for example the type of sort u is sort (succ u). The one notable case here is param \$i\$, which represents the $i$th universe variable, where $i<n$ where $n$ is the number of universe parameters to the current declaration.
This definition corresponds to the following BNF style grammar from (Carneiro, 2019), with the main change being that it is more precise about how variables and binding are represented. (For presentational reasons, we will use named variables in the informal version, but the formalization uses lift and subst functions when moving expressions between contexts.)
$\ell$ $::=0\mid\operatorname{S}\ell\mid\max(\ell,\ell)\mid\operatorname{imax}(\ell,\ell)\mid u$
@@
2.2. Typing and definitional equality
@@
Figure 1. The rules for the judgment $\Gamma\vdash e\equiv e^{\prime}:\alpha$, with parameters $E$ (the global environment) and $n$ (the number of universe parameters in context), suppressed in the notation. This is VEnv.IsDefEq in the formalization.
For the typing rules, we use a slight variation of the typing judgments from (Carneiro, 2019), see Figure 1. Some remarks:
  • In (Carneiro, 2019), there were two judgments $\Gamma\vdash e:\alpha$ and $\Gamma\vdash e_{1}\equiv e_{2}$ which are mutually inductive (due to t-conv and t-proof-irrel), and $\Gamma\vdash e_{1}\equiv e_{2}:\alpha$ was a defined notion for $\Gamma\vdash e_{1}\equiv e_{2}\ \wedge\ \Gamma\vdash e_{1}:\alpha\ \wedge\ \Gamma\vdash e_{2}:\alpha$. In this version, we have only one all-in-one relation $\Gamma\vdash e_{1}\equiv e_{2}:\alpha$, and we define $\Gamma\vdash e:\alpha$ to mean $\Gamma\vdash e\equiv e:\alpha$ and define $\Gamma\vdash e_{1}\equiv e_{2}$ as $\exists\alpha.\;(\Gamma\vdash e_{1}\equiv e_{2}:\alpha)$.
    We conjecture the two formulations to be equivalent, but this version seems to be easier to prove basic structural properties about (see subsection 2.3). Moreover Lean does not have good support for mutual inductive predicates, so keeping it single-inductive makes induction proofs easier.
@@
  • The t-extra rule says that we can add arbitrary definitional equalities from the environment. This is how we will add the two supported “extensions” to the theory, for inductive types and quotients, without the details of these extensions complicating reasoning about the core theory. But we can’t actually support arbitrary definitional equalities—instead each theorem about the typing judgment has its own assumptions about what extensions are allowed. (However, note that the specific constraints on t-extra are still unfinished, pending future work on inductive type specification.)
2.3. Properties of the typing judgment
@@
  1. (1)
    (IsDefEq.closedN) If $\Gamma\vdash e\equiv e^{\prime}:\alpha$ and $\Gamma$ is closed, then $e$, $e^{\prime}$, and $\alpha$ are closed (all free variables have indices less than $|\Gamma|$).
  2. (2)
@@
  3. (3)
    (IsDefEq.instL) If $\Gamma\vdash_{E,n}e\equiv e^{\prime}:\alpha$ and $\forall i.\;n^{\prime}\vdash\ell_{i}\ \mathsf{ok}$, then $\Gamma[\bar{u}\mapsto\bar{\ell}]\vdash_{E,n^{\prime}}e[\bar{u}\mapsto\bar{\ell}]\equiv e^{\prime}[\bar{u}\mapsto\bar{\ell}]:\alpha[\bar{u}\mapsto\bar{\ell}]$.
  4. (4)
    (IsDefEq.instN) If $\Gamma,x:\beta\vdash e_{1}\equiv e_{2}:\alpha$ and $\Gamma\vdash e_{0}:\beta$, then $\Gamma\vdash e_{1}[x\mapsto e_{0}]\equiv e_{2}[x\mapsto e_{0}]:\alpha[x\mapsto e_{0}]$.
Proof sketch.
  1. (1)
    This is a straightforward proof by induction on $\Gamma\vdash e\equiv e^{\prime}:\alpha$, except that one gets stuck at t-const because (assuming the environment is well-typed, see section 4) we know that $\vdash c_{\bar{\ell}}:\alpha$ and need that $\alpha$ is closed. So in fact this is a double induction, first over the environment $E$ and then over $\Gamma\vdash e\equiv e^{\prime}:\alpha$. We also need lemmas about $e[x\mapsto e^{\prime}]$ and $e[\bar{u}\mapsto\bar{\ell}]$ preserving closedness, but these are also direct by induction on $e$.
  2. (2)
    By induction. We use Lemma 2.1.1 in the t-const case, to show that the $\alpha$ in $\Gamma,\Gamma^{\prime}\vdash c_{\bar{\ell}}\equiv c_{\bar{\ell}^{\prime}}:\alpha$ can be used in $\Gamma,\Delta,\Gamma^{\prime}\vdash c_{\bar{\ell}}\equiv c_{\bar{\ell}^{\prime}}:\alpha$ without renaming any bound variables because $\alpha$ is closed.
  3. (3)
@@
  4. (4)
    By induction on the first hypothesis. This uses
    Lemma 2.1.2 when lifting $\Gamma\vdash e_{0}:\beta$ under binders.
@@
Proof.
We have $\Gamma,x:\alpha^{\prime}\vdash x:\alpha$ by t-bvar, t-conv and weakening, and $\Gamma,\_:\alpha^{\prime},x:\alpha\vdash e_{1}\equiv e_{2}:\beta$ by weakening, so $\Gamma,x:\alpha^{\prime}\vdash e_{1}[x\mapsto x]\equiv e_{2}[x\mapsto x]:\beta[x\mapsto x]$ by Lemma 2.1.4. ∎
Lemma 2.3 (HasType.forallE_inv).
If $\Gamma\vdash(\forall x:\alpha.\;\beta):\gamma$ then $\Gamma\vdash\alpha\ \mathsf{type}$ and $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$ (and therefore also $\Gamma\vdash(\forall x:\alpha.\;\beta)\ \mathsf{type}$).
Proof sketch.
@@
  • t-all: If $e_{1}$ is $\forall x:\alpha.\;\beta$ we are done; if $e_{2}$ is $\forall x:\alpha.\;\beta$ then we obtain $\Gamma,x:\alpha^{\prime}\vdash\beta\ \mathsf{type}$ and need Lemma 2.2 to get $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$.
  • t-beta: For this to apply, it must be that we have $(\lambda y:\delta.\;e)\;e^{\prime}\equiv e[y\mapsto e^{\prime}]$ where $e[y\mapsto e^{\prime}]$ is $\forall x:\alpha.\;\beta$. So either $e=y$ and $e^{\prime}=\forall x:\alpha.\;\beta$, in which case the inductive hypothesis for $e^{\prime}$ applies, or $e=\forall x:\alpha^{\prime}.\;\beta^{\prime}$ with $\alpha^{\prime}[y\mapsto e^{\prime}]=\alpha$ and $\beta^{\prime}[y\mapsto e^{\prime}]=\beta$ in which case $\Gamma,y:\delta\vdash\alpha^{\prime}\ \mathsf{type}$ and $\Gamma,y:\delta,x:\alpha^{\prime}\vdash\beta^{\prime}\ \mathsf{type}$ by the inductive hypothesis for $e$, and Lemma 2.1.4 applies.
@@
Lemma 2.4 (HasType.sort_inv).
If $\Gamma\vdash_{E,n}\mathsf{U}_{\ell}:\gamma$ then $n\vdash\ell\ \mathsf{ok}$ (and therefore also $\Gamma\vdash\mathsf{U}_{\ell}:\mathsf{U}_{\operatorname{S}\ell}$).
Proof sketch.
Similar to Lemma 2.3. ∎
@@
If $\vdash\Gamma\ \mathsf{ok}$ and $\Gamma\vdash e:\alpha$ then $\Gamma\vdash\alpha\ \mathsf{type}$.
Proof sketch.
@@
  • t-app: We have $\Gamma\vdash(\forall x:\alpha.\;\beta)\ \mathsf{type}$ from the IH and $\Gamma\vdash e_{2}:\alpha$ by assumption, and from Lemma 2.3 we get $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$, so $\Gamma\vdash\beta[x\mapsto e_{2}]\ \mathsf{type}$ by the substitution lemma.
  • t-all: We have $\Gamma\vdash\mathsf{U}_{\ell_{1}}\ \mathsf{type}$ and $\Gamma,x:\alpha\vdash\mathsf{U}_{\ell_{1}}\ \mathsf{type}$ from the IH, so by Lemma 2.4 we have $n\vdash\ell_{1},\ell_{2}\ \mathsf{ok}$, therefore $\Gamma\vdash\mathsf{U}_{\operatorname{imax}(\ell_{1},\ell_{2})}\ \mathsf{type}$.
@@
If $\Gamma,x:\alpha\vdash f\equiv f^{\prime}:\beta$ and $\Gamma\vdash a\equiv a^{\prime}:\alpha$ then $\Gamma\vdash f[x\mapsto a]\equiv f^{\prime}[x\mapsto a^{\prime}]:\beta[x\mapsto a]$.
Proof sketch.
This is a generalization of Lemma 2.1.4, but we need Theorem 2.5 to prove that $\alpha$ and $\beta$ are well-typed to apply the rules below.
@@
$f[x\mapsto a]$ $\equiv(\lambda x:\alpha.\;f)\;a$
@@
To finish, we apply the lemma twice, once with $f$ and $\beta$ so that it suffices to show $\beta[x\mapsto a]\equiv\beta[x\mapsto a^{\prime}]:\mathsf{U}_{\ell}$ and then again with $\beta$ in place of $f$ and $\mathsf{U}_{\ell}$ in place of $\beta$ so that it suffices to show $\mathsf{U}_{\ell}[x\mapsto a]\equiv\mathsf{U}_{\ell}[x\mapsto a^{\prime}]:\mathsf{U}_{\operatorname{S}\ell}$, which is true by reflexivity. ∎
2.4. Conjectured properties of the typing judgment